OpenStudy (anonymous):

What is the solution of the equation? 4p/6 + 27= 39 A. 22 B. 99 C. 4 D. 18

OpenStudy (anonymous):

We want to isolate p so the first step would be to subtract 27 to both sides. What do you get then?

OpenStudy (anonymous):

can you please explain a little more

OpenStudy (anonymous):

We want p to be by itself on one side. That means that everything expect the p on the left hand side must go away. We have a positive 27 on the left hand side. We want to get rid of it and we do this by subtract it by 27. But we must remember if we do something on one side we must also do the same exact thing on the other side. That's why I said subtract both sides by 27. Understand?

OpenStudy (anonymous):

ok i got it i think but how would i subtract 27 from a fraction

OpenStudy (anonymous):

you don't it cancels out the +27 and it goes away.

OpenStudy (anonymous):

scratches head umm ok so 27 minus 27 equals zero

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4p/6 + 27= 39 Subtract 27 to both sides 4p/6 + 27-27= 39-27 now you get 4p/6 + 0= 39-27 then the right sides simplifies to 4p/6 + 0= 12 zero goes away 4p/6 = 12

OpenStudy (anonymous):

Follow so far?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

It's ok if you get confuse just read over the posts again. We have this now 4p/6 = 12 I'm going write it in a program so you can see it in a different way. It's still the same problem it just looks nicer now \[\frac{4}{6}p= 12 \]

OpenStudy (anonymous):

You then have to multiply both sides by 6. This will get rid of the 6 in the fraction from the left hand side.

OpenStudy (anonymous):

You then have to multiply both sides by 6. This will get rid of the 6 in the fraction from the left hand side. \[6*\frac{4}{6}p =12*6\\4p =12*6\\4p=72\]

OpenStudy (anonymous):

The last step will be to divide both sides by 4. This will get rid of the 4 on the left hand side.

OpenStudy (anonymous):

18

OpenStudy (anonymous):

yes you are right p=18 good job!

OpenStudy (anonymous):

dude you rock lol feel like helping me with a few more??